OpenStudy (anonymous):

A hot bowl of soup cools according to Newton's law of cooling. Its temperature (in degrees Fahrenheit) at time t is given by the function t(t)=68+144e^-0.04t , where t is expressed in minutes. At what time will the soup reach 200oF? A. 137.39 minutes B. 2.18 minutes C. 16.39 minutes D. 1.02 minutes

OpenStudy (anonymous):

Set the equation equal to 200 then subtract 200 from each side and solve the quadratic equation. I'm going to assume that you need the quadratic formula to continue.

OpenStudy (anonymous):

problem is idk how to solve it

OpenStudy (campbell_st):

ok.... you should set t(t) = 200 so \[200 = 68 + 144e^{-0.04t}\] set 1 subtract 68 from both sides.... what do you get..?

OpenStudy (anonymous):

-132+144e^-0.04t

OpenStudy (campbell_st):

nearly \[132 = 144e^{-0.04t}\] you are solving for t like any other equation, so you want t on one side and the value on the other side... next step divide both sides by 144

OpenStudy (anonymous):

.916=e^-0.04t

OpenStudy (campbell_st):

ok.... so do you know log laws...? particularly \[\ln{e^a} = a\]

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i have my calculator though

OpenStudy (campbell_st):

ok.... well you'll have to trust me... do you have a calculator with an ln function

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

pk this is what you need to find ln(132/144) it will be a negative answer...

OpenStudy (anonymous):

-.087011377

OpenStudy (campbell_st):

i think you're missing a zero after the decimal point... but thats ok by taking the ln on both sides the exponential disappears... so ln(132/144) = - 0.04t or -0.087011 = -0.04t last step is to divide both sides by -0.04

OpenStudy (anonymous):

2.175284425

OpenStudy (campbell_st):

great... that is how many minutes it takes the soup to cool from 212 to 200 degrees... when t = 0 the temperature of the soup is \[t(t) = 68 + 144e^{-0.04 \times 0} = 68 + 144 = 212\]

OpenStudy (campbell_st):

so 2.1752... minutes later the soup is 200 degree

OpenStudy (campbell_st):

hope it makes sense...